OpenStudy (anonymous):

Nancy bought 360 tropical fish for a museum display. She bought 8 times as many goldfish as angelfish. How many of each type of fish did she buy? Which is a system of equations to model the problem if x represents the number of angelfish Nancy bought, and y represents the number of goldfish she bought? x – y = 360 y = 8x x + y = 360 y = 8x x + y = 8 y = 360x x + 8y = 360 y = 8x

OpenStudy (solomonzelman):

Goldfish=g Angelfish=a g=a * 8 a+g=360 a * 8 + a = 360 9a=360 a=40 So she bought 40 of angelfish and knowing that goldfish is 8 times more 40 * 8 which is equal to 320 so she bought 320 of goldfish

OpenStudy (amoodarya):

Goldfish=x Angelfish=y as "SolomonZelman" says x+y=360 y=8x

OpenStudy (anonymous):

thank u to the both of u for explaining

OpenStudy (solomonzelman):

You welcome!